OpenStudy (anonymous):

find the following limit. lim t->0 (1/(t -1)/(t^2+t))

OpenStudy (anonymous):

sorry is \[\lim_{t \rightarrow 0} (1/t-1/t^2+t)\]

OpenStudy (anonymous):

if uses the lcd i get t^2/t^3+t^2

OpenStudy (anonymous):

is this the expression? \[\huge \frac{\frac{1}{t-1}}{t^2+t} \]

OpenStudy (anonymous):

\[\huge \lim_{t\rightarrow 0}\frac{\frac{1}{t-1}}{t^2+t} \]

OpenStudy (anonymous):

no is minus (1/t-1)- 1/t^2+t

OpenStudy (zepp):